convert Sn=(6+66+666....Nth term)
into GP and find Nth term

Grade:12th Pass

4 Answers

Aman Bansal
11 years ago

Dear Rohan,

The given sequence is 8, 88, 888, 8888…
This sequence is not a G.P. However, it can be changed to G.P. by writing the terms as
Sn = 8 + 88 + 888 + 8888 + …………….. to n terms

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

shubham tibra
11 years ago

Sn =6/9(9+99+999....n)     sn= 6/9(10+100+1000...n -1-1-1-1.....n) sn =6/9(gp with a=10 and r=10 upto nterm - summation of 1 i.e. n) solving it u will get the answer

anshumanpandey
7 years ago
By anshuman pandey AIS With arpit and ankit pandey
Its too easy (6+66+666+6666+.............n terms)
 
take common 6 ( 1+11+111............)  multiply by 9
then 6/9 (9+99+999+.........)
 6/9(10+100+1000+......-1-1-1-1 n)
Helper
6 years ago
6+66+666+..6(1+11+111+...)Multiply&Divide by 96/9(9+99+999+.....)6/9[(10-1)+(100-1)+(1000-1)+...]6/9[10+10²+10³+....]-[(1)(1)...... n times]Here a=10 and r=10 Hence, S = a.r^n-1/ r-1S= 10(10^n-1)/10-1S=10(10^n-1)/9=> 6/9[10/9(10^n-1) - n]